OpenStudy (victoriaf16):

Can someone help me solve this? A = 350(1 + 0.06/12)^(12*2)

OpenStudy (photon336):

\[350(1 + \frac{ 0.06 }{ 12 })^{12*2} \]

OpenStudy (photon336):

@Victoriaf16 this looks like compound interest formula

OpenStudy (victoriaf16):

I think thats it. but it doesnt make sense to me to divide 0.06 by 12.

OpenStudy (photon336):

well, remember PEMDAS? we've got to work with what's in the ( ) first

OpenStudy (photon336):

if this is compound interest the formula looks something like this. \[A = A_{0}(1+\frac{ a }{ n })^{nt}\] 0.06/12 = 0.005 we take care of that that first. then add that to 1 to get 1.005 now we've got \[360(1.005)^{24}\] P E M D A S now we raise that number to the 24th power then multiply by 260

OpenStudy (victoriaf16):

you were right about it being a compound interest formula

OpenStudy (victoriaf16):

where did you get 360?

OpenStudy (photon336):

mistake that was supposed to be 350

OpenStudy (victoriaf16):

oh okay

OpenStudy (agent0smith):

@Victoriaf16 is your question how to work this out, or why the 0.06 is being divided by 12 in the first place? It's probably because the interest is compounded monthly, and there's 12 months in a year, so you need to divide the interest rate by 12.

OpenStudy (photon336):

yeah and the duration is 2 years

OpenStudy (victoriaf16):

I mean, theyre both my question. My original question is the answer to this.

OpenStudy (photon336):

so you want to figure out how much money you have when the interests is compounded monthly at a rate of 0.06% after 2 years.

OpenStudy (agent0smith):

In that case it'd be better to post the actual question, ie what you were given to start with

OpenStudy (photon336):

well, remember we can't give the direct answer, try to post the original question with all the information for us.

OpenStudy (victoriaf16):

Larry and Peggy are making decisions about their bank accounts. Larry wants to deposit $350 as a principle amount, with an interest of 4% compounded quarterly. Peggy wants to deposit $350 as the principle amount, with an interest of 6% compounded monthly. Explain which method results in more money after 2 years. Show all work. this is the original question.

OpenStudy (victoriaf16):

i somehow got to this equation yet i have no idea how to solve it.

OpenStudy (agent0smith):

You just have to plug numbers into the compound interest formula \[\LARGE A = P\left( 1+\frac{ r }{ n } \right)^{n t}\]

OpenStudy (agent0smith):

P is the initial deposit, r is interest rate, n is number of times compounded per year, t is years

OpenStudy (victoriaf16):

Well thats what i did, but idk how to solve after that.

OpenStudy (agent0smith):

But there's Larry and Peggy, you only did one of them. What's so hard about solving it? Just plug it into a calculator...

OpenStudy (agent0smith):

\[\Large 350(1 + \frac{ 0.06 }{ 12 })^{12*2}\]You can simplify some of it by hand, but you can also just plug it into a calculator, as long as you type it in correctly.

OpenStudy (victoriaf16):

well then im not typing it in correctly cause what im getting makes no sense

OpenStudy (agent0smith):

Then simplify some of it, bit by bit. First work out 1+0.06/12 (type that into a calculator if you need to). And the 2*12.

OpenStudy (victoriaf16):

I know how to solve them apart but how do i put them together? @agent0smith

OpenStudy (agent0smith):

What do you mean?

OpenStudy (victoriaf16):

i got 1.005 and 24. Am i adding those or multiplying them? How am i putting it with the 350? It just doesnt make sense as a whole to me.

OpenStudy (agent0smith):

350*(1+0.06/12)^24 ^ means exponent

OpenStudy (victoriaf16):

that would give me an insanely high amount

OpenStudy (agent0smith):

No it won't. Just work it out as i just wrote it, type it into a calculator.

OpenStudy (victoriaf16):

Well im trying but its definitely giving me an insane amount no matter how i do it

OpenStudy (agent0smith):

You can literally copy and paste 350*(1+0.06/12)^24 into google and get the correct answer. If you're using a calculator, i'd suggest doing (1+0.06/12)^24 first, and telling me what you get. Then multiply that by 350

OpenStudy (agent0smith):

I'm assuming you're not using parentheses exactly as they're shown, if you are getting a huge number.

OpenStudy (victoriaf16):

Im getting 1.005 from whats in the parentheses, is that wrong or?

OpenStudy (agent0smith):

That's fine, now do 350*(1.005)^24

OpenStudy (victoriaf16):

i got 394.505 from google

OpenStudy (agent0smith):

But you need to be able to do it on a calculator too. Do 350*(1.005)^24 EXACTLY as it's written.

OpenStudy (victoriaf16):

well then i have to multiply 1.005 by itself 24 times, im only human

OpenStudy (agent0smith):

Uh, no... you don't do that by hand unless you're insane 350*(1.005)^24 on a calculator.

OpenStudy (agent0smith):

* means multiply and ^ means exponent

OpenStudy (campbell_st):

the thing nobody seems to have explained, lookking at your formula is that the compounding period are monthly... so if the interest is 6% p.a. then per month the interest is 6/12 = 0.5% now change that to a decimal. the investment is for 2 years... so if it is compounded monthly 12 x 2 = 24 is the value of n the number of compounding periods... hope it helps